OpenStudy (anonymous):

if 2^x = 15^y = 3^z = 10^w yhen 1/x + 1/y :1/z+ 1/w is?

OpenStudy (experimentx):

not sure this will work let, 2^x = 15^y = 3^z = 10^w = k x = k/ln2

OpenStudy (experimentx):

sorry, lnk/ln2

OpenStudy (experimentx):

Oops ... we can express it interms of z

OpenStudy (anonymous):

@experimentX u r 1st way is correct plzz go on

OpenStudy (experimentx):

\[ \frac{\ln 2+\ln 15 + \ln 10}{\ln k} = \frac{\ln 300}{\ln k}\]

OpenStudy (anonymous):

wat is in

OpenStudy (experimentx):

I think we have to replace k with z

OpenStudy (anonymous):

ok then

OpenStudy (experimentx):

\ln k = z \ln 3 i guess \[ \frac{\ln 300}{z\ln3}\]

OpenStudy (anonymous):

@experimentX any other way!

OpenStudy (experimentx):

lol ... i'm not quite good with algebra ... if you know the answer, i might able to connect it.

OpenStudy (anonymous):

sorry for the mistake in question plz look the question now

OpenStudy (anonymous):

if 2^x = 15^y = 3^z = 10^w yhen 1/x + 1/y :1/z+ 1/w is?

OpenStudy (experimentx):

Oh ... sorry, i didn't see the : in the question

OpenStudy (anonymous):

@experimentX it was not there i just edited now

OpenStudy (experimentx):

use same approach ... cancel out ln k

OpenStudy (anonymous):

sorry for the mistake

OpenStudy (experimentx):

\[ \frac{\ln 30}{\ln 300}\]

OpenStudy (anonymous):

The answer should be 1:1 @experimentX

OpenStudy (experimentx):

OH ... let's see where i made mistake

OpenStudy (experimentx):

Oh .. sorry again, i thought that 3 was 30

OpenStudy (anonymous):

i did nt understand those things plz do in other way

OpenStudy (anonymous):

@FoolForMath

OpenStudy (experimentx):